### v2

# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

Abstract
Purpose: This paper introduces a novel graph-based method, GARG-AML, for efficient and effective anti-money laundering (AML). It quantifies smurfing risk, a popular money laundering method, by providing each node in the network with a single interpretable score. The proposed method strikes a balance among computational efficiency, detection power and transparency. Different versions of GARG-AML are introduced for undirected and directed networks.
Methodology: GARG-AML constructs the adjacency matrix of a node’s second-order neighbourhood in a specific way. This allows us to use the density of different blocks in the adjacency matrix to express the neighbourhood’s resemblance to a pure smurfing pattern. GARG-AML is extended using a decision tree and gradient-boosting classifier to increase its performance even more. The methods are tested on synthetic and on open-source data against the current state-of-the-art in AML.
Findings: We find that GARG-AML obtains state-of-the-art performance on all datasets. We illustrate that GARG-AML scales well to massive transactions graphs encountered at financial institutions. By leveraging only the adjacency matrix of the second-order neighbourhood and basic network features, this work highlights the potential of fundamental network properties towards advancing fraud detection.
Originality: This paper uses only basic network features and expert knowledge on smurfing to construct a performant AML system. The originality lies in the translation of smurfing detection to these features and network representation. Our proposed method is built around the real business needs of scalability and interpretability. It therefore provides a solution that can be easily implemented at financial institutions or incorporated in existing AML solutions.
Keywords Fraud Analytics $\cdot$ Anti-Money Laundering $\cdot$ Network Analytics $\cdot$ Smurfing $\cdot$ Microstructuring
1 Introduction
Money laundering disguises illegally-obtained funds, often from criminal activities, to make them appear legitimate. Once cleaned, the money can be used to finance further crimes. Estimates put the amount of laundered money at $2\%$ to $5\%$ of global GDP, which amounts to USD 2 trillion annually (United Nations Office on Drugs and Crime, (n.d.). Since laundering money involves interactions with and through the normal financial system, regulators have imposed stringent anti-money laundering (AML) rules on financial institutions (FATF, 2012-2023; Levi & Reuter, 2006).
Money laundering involves three main steps (United Nations Office on Drugs and Crime, (n.d.; Levi & Reuter, 2006). Placement refers to the introduction of funds into the financial system. During layering, money is moved around to obscure its illegal origins. Finally, integration is the incorporation of laundered money into broader circulation. Because of the involvement of different steps and actors, network analytics has become an important tool for AML (Deprez et al., 2024).
A popular money laundering method is smurfing (Madinger, 2012; Welling, 1989; Starnini et al., 2021; Zhdanova et al., 2014). Large transactions between two accounts are often flagged according to business rules (Bolton & Hand, 2002; Senator et al., 1995; Oztas et al., 2024) in accordance with regulation (Welling, 1989). Therefore, transactions are split into smaller amounts and transferred using multiple money mules—also called smurfs—making detection of illicit money flows harder since (1) the origin of the funds is obscured and (2) smaller transactions seem less suspicious.
Two types of patterns are typical for smurfing (Starnini et al., 2021): (1) scatter-gather and (2) gather-scatter (Altman et al., 2023). Both patterns are visualised in Figure 1. With scatter-gather, money is sent to its destination account via money mules. Gather-scatter, on the other hand, consists of collecting money from multiple accounts into one (or a few) account(s) and sending it to many destination accounts. Gather-scatter is a popular money laundering method for cryptocurrencies (Wu, Hu, Zhou, Wang, Luo, Wang, Zhang & Ren, 2021).
Figure 1: The two smurfing patterns described in the literature (Starnini et al., 2021; Altman et al., 2023); as scatter-gather (left) and gather-scatter (right).
To fight money laundering, AML practices use customer and transaction data to estimate each client’s money laundering propensity. When suspicion is raised, investigators build a suspicious activity report (SAR) to file with the regulator (FATF, 2012-2023; Levi & Reuter, 2006). Given the vast volume of monitored transactions and the involvement of different business departments in filing a SAR, two aspects are key to ensuring the effective business adoption of a method:
  • The method must process large volumes of transactions in a memory and time efficient manner;
  • The predictions must be interpretable for smooth integration in SAR filing.
Despite these business requirements, the current AML literature often focuses on deep graph neural networks (GNNs) (Deprez et al., 2024). GNNs are, however, not interpretable, prohibiting the adoption of these novel network analytics methods in practice. Consequently, many institutions still rely on business rules and manual feature engineering based on expert knowledge, which are expensive to maintain and easy to evade.
Given the specific nature of smurfing and the operational requirements faced by financial institutions, we introduce Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), a novel method that grades clients on the basis of their second-order neighbourhood in the transaction network. GARG-AML operates by analysing the resulting adjacency matrix and computing scores based on the contrast between high-density and low-density regions to quantify the network’s resemblance to smurfing. We introduce two versions of GARG-AMM, one based on directed and one on undirected transaction networks.
Moreover, we present an extension to the base GARG-AML score, increasing smurf detection power by combining GARG-AML with classic network centrality measures and tree-based learners. This shows the versatility and power of our method, without compromising on the business needs of speed and simplicity.
The performance of GARG-AML and its extensions is evaluated against state-of-the-art smurfing detection methods using both synthetic and two open-source network datasets. These datasets, which include millions of clients and transactions, closely resemble real-world financial networks. As a result, our experiments assess both detection accuracy and computational efficiency.
The main contributions of this work are as follows:
  • We introduce a novel smurfing detection method that is scalable to large transaction networks, interpretable and easily integrated into existing machine learning models;
  • We evaluate our method on multiple datasets which are uniquely suited for assessing the effectiveness of our approach;
  • We compare the current state-of-the-art smurfing detection methods, and we perform the first analysis of the scalability of these methods to very large transaction datasets;
  • Complete implementations of the methods and experiments are provided on GitHub to promote further adoption of and experimentation on the GARG-AML scores by peer researchers and practitioners.
The remainder of this work is organised as follows. Related work on anti-money laundering and smurfing is discussed in Section 2. The methodology behind the GARG-AML score calculation is described in Section 3. Section 4 outlines the experimental evaluation, and the results are presented in Section 5. We discuss the main limitations, in terms of both the data in general and GARG-AML in particular, in Section 6. Section 7 concludes this work.
2 Related Work
Anti-money laundering (AML), like other areas of financial fraud detection, traditionally relied on IF-THEN business rules—combinations of expert knowledge and regulatory requirements (Bolton & Hand, 2002; Leonard, 1995; Sánchez et al., 2009; Welling, 1989; Cardoso et al., 2022; Deprez et al., 2024). These rules typically monitor individual transactions or aggregated amounts over time (FATF, 2012-2023). While offering transparency in decision-making (Zhu & Hu, 2013), they are costly to maintain (Van Belle et al., 2023) and can be easily circumvented once exposed. As a result, machine learning has emerged as a preferred approach, offering scalability and the ability to detect complex laundering patterns (Chen et al., 2018).
To capture interactions among people, network analytics for AML has been studied extensively in recent years (Deprez et al., 2024). Research has often relied on manual network feature engineering for profile definition. Dreżewski et al. (2015) used different centrality measures, which resulted in nine roles in a criminal network. Fronzetti Colladon & Remondi (2017) used similar metrics across industries and analysed the correlation with the target label to demonstrate the added value of network analytics for AML. Wu, Liu, Chen, Huang, Zheng & Zhang (2021) define temporal motives specifically tailored to money laundering and mixing detection in the Bitcoin network.
In addition to manual feature engineering, graph neural networks (GNNs) have also been studied extensively (Lo et al., 2023; Alarab et al., 2020; Weber et al., 2019; Deprez et al., 2024). These methods capture the complex geometry of suspicious money laundering activities in financial networks by learning network topology while also incorporating node features. A drawback is that GNNs are not interpretable (Kute et al., 2021), which limits their use by investigators when filing their reports.
Smurfing has long been a popular laundering method (Welling, 1989; Senator et al., 1995). Welling (1989) suggested that it has been applied since 1970, when the Bank Secrecy Act was passed in the USA, which required reporting of transactions above $\mathdollar 10,000$. To this day, similar rule-based systems are in place that require immediate investigation of transactions that surpass a given amount (Senator et al., 1995; Oztas et al., 2024; Bolton & Hand, 2002).
Research has already been conducted on smurfing detection. Zhdanova et al. (2014) introduced PSA@R for detecting smurfing in mobile money networks. It uses different underlying workflow models. Transaction streams that differ from the process model are flagged as anomalous. Additional criteria are applied to find money mules on the basis of the difference in the amounts of money coming into and going out of an account.
FlowScope (Li et al., 2020) tries to find anomalous sub-graphs on the basis of money flow. The authors constructed a k-partite network with two sets of nodes: nodes with net inflow to or net outflow from the bank and the remaining nodes, which correspond to the middle men. Li et al. (2020) defined their anomaly score on the basis of the difference between the inflow and outflow of money at these middle men, and they claimed that smurfs transfer almost all the money they receive.
A similar idea was put forward with AutoAudit (Lee et al., 2020). The full adjacency matrix of the network is reordered, and depending on the density of the different blocks, a group of transactions is flagged as smurfing.
Starnini et al. (2021) presented a method that merges two transaction datasets and retains only those connections where both transactions occur within a specific time window. Prior to this merger, a substantial amount of data, i.e., nodes deemed trustworthy by the bank together with a majority of transactions, is discarded. After merging, the method retains certain motifs and applies an additional post-processing step, which then triggers manual investigations by AML experts.
SMoTeF (Shadrooh & Nørvåg, 2024) is a three-step approach. The first is smurfing pattern extraction, using existing methods. The novelty comes from the second and third steps, where the patterns are checked against temporal constraints and maximal money flow, respectively. Many transactions are again pruned before making final predictions.
[…]
Closely related to smurfing is mixing, which is a money laundering method that is used for cryptocurrencies (Liu et al., 2021). Coins are transferred to the same entity and are redistributed afterward (Wu, Hu, Zhou, Wang, Luo, Wang, Zhang & Ren, 2021). This obfuscates the history of ownership, which results in an inability to know the money’s real origin. Mixing can be viewed as the second smurfing pattern in Figure 1, i.e., gather-scatter.
An anomaly score for mixing detection was introduced by Prado-Romero et al. (2018). The authors claimed that mixing services act as bridges between communities in a transaction network by linking otherwise unrelated users. After community detection, the anomaly score expresses the difference in a node’s intercommunity links compared with those of the other nodes in the same community.
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) constructed two networks for network feature extraction. The first links accounts using their transactions, whereas the second network represents both accounts and transactions as nodes to leverage transaction information. The extracted network features are based on predefined temporal motifs, which are added to the account and transaction features to detect smurfing in a positive and unlabelled (PU)-learning setting.
The above literature is summarised and juxtaposed against our proposed method according to the methodology in Table 1 and according to the experimental set-up in Table 2. As expected, most methods for anti-money laundering on fiat currency consider only scatter-gather patterns, while research on crypto-currency is mostly concerned with gather-scatter.
The evaluation in most papers is done on relatively small datasets, making it harder to judge whether they scale well to the massive transaction networks faced by financial institutions. The dataset that was used in most of the above papers is the Czech dataset (Data World, 1999), which is a rather small network that contains approximately 11,000 nodes and approximately 273,000 edges. This dataset is unlabelled, so the authors of each paper inject synthetic pattern. Additionally, in most papers, evaluation and comparison is limited to threshold-dependent metrics like precision, recall and F1.
Table 1: Summary of key aspects of the methodology in the literature.
Paper Pattern Client Scoring Parallel Calculation All Nodes Code Scalable
Zhdanova et al. (2014) Scatter-gather
Li et al. (2020) Scatter-gather
Lee et al. (2020) Scatter-gather
Starnini et al. (2021) Scatter-gather & gather-scatter
Shadrooh & Nørvåg (2024) Scatter-gather
Prado-Romero et al. (2018) Gather-scatter
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) Gather-scatter
This work Scatter-gather & gather-scatter
Table 2: Summary of key aspects of the experimental set-ups in the literature
Paper Dataset Size Dataset (nodes) Metrics
Zhdanova et al. (2014) Simulated 10,000 Precision, recall
Li et al. (2020) Proprietary & Czech Financial 6,130,000 & 11,374 F1, Time
Lee et al. (2020) Czech Financial & Enron 11,374 & 16,771 Accuracy
Starnini et al. (2021) Proprietary 18,000,000 (reduced to 21,000) Time
Shadrooh & Nørvåg (2024) Czech Financial & Enron & Darpa 11,374 & 16,771 & 371,000 Time, Precision, Recall, F1
Prado-Romero et al. (2018) Bitcoin 942,204 Recall
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) WalletExplorer $\pm 2,500,000$ TPR, FPR, geometric mean
This work Simulated & IBM AML 100 & 10,000 & 100,000 & 515,080 & 2,054,390 Time, AUC-ROC, AUC-PR, Rank
In this paper, we propose a new method that scores all clients individually, where score can be calculated in parallel. These score can also be used both for scatter-gather as well as for gather-scatter patterns. The experiments are done on very large, realistic data, where we use threshold-independent metrics that allow us to have a comprehensive comparison among the methods under consideration.
3 GARG-AML
The main idea of GARG-AML is to provide each individual node with a score expressing how likely it is that this entity is involved in a smurfing scheme. This is done by analysing key characteristics of the second-order neighbourhood of the nodes. Below, we start by formally defining what we mean with second-order neighbourhood. Afterwards, we provide two version for GARG-AML, one for undirected and one for directed networks. The full GARG-AML method is given in Algorithm 1 in the experimental set-up in Section 4.2.
3.1 Preliminaries and Notation
A network $G(V,E)$ is defined by a set of nodes, $V=\{v_{1},\ldots,v_{k}\}$, and a set of edges, $E\subset V\times V$. The edges can be represented using the adjacency matrix, $A$, where the elements are defined as
$A_{ij}=\begin{cases}1&\text{if }(v_{i},v_{j})\in E\\ 0&\text{elsewhere}\end{cases}$
Suppose $\mathcal{N}_{k}(v)$ represents the set of nodes that are at distance exactly $k$ from node $v$, with $\mathcal{N}_{0}(v)=\{v\}$. For GARG-AML, we consider the neighbours of each node $v$ at distances one and two, i.e., $\mathcal{N}_{1}(v)$ and $\mathcal{N}_{2}(v)$, respectively. Suppose that $\mid\mathcal{N}_{1}(v)\mid=n$ and $\mid\mathcal{N}_{2}(v)\mid=m-1$; then, the adjacency matrix of the second-order neighbourhood of $v$ is of size $(m+n)\times(m+n)$.
The $k$th-order neighbourhood of node $v$ is the sub-network $G(V^{\prime},E^{\prime})$ where $V^{\prime}=\bigcup\limits_{l\leq k}\mathcal{N}_{k}(v)$ and $E^{\prime}=\{(u,w)\mid u,w\in V^{\prime}\vee(u,w)\in E\}$. Hence, it is the sub-network that representing all nodes at distance at most $k$ and all the interaction among these nodes.
3.2 Undirected
We introduce GARG-AML using the smurfing pattern defined by Starnini et al. (2021) and illustrated in Figure 2. In this example, we want to give node $A$ a score. The adjacency matrix of its second-order neighbourhood, given in Equation (1), is constructed by arranging the nodes in a specific order: node $A$ is placed first, followed by the second-order neighbour $E$, and then its first-order neighbours $B,C$ and $D$. This ordering results in an adjacency matrix that is divided into four distinct blocks. The diagonal blocks contain only zeros, whereas the off-diagonal blocks contain ones.
In this example, we will consider only the scatter-gather pattern, as this is most widely recognised as smurfing. The GARG-AML calculations and node order stay the same in the gather-scatter case.
Figure 2: Example of a smurfing pattern. AA is the node that will receive the score. The first-order neighbours ($B,C,D$) are shown in purple, and the second-order neighbour ($E$) is shown in green.
$\begin{array}[]{r}A\\ E\\ B\\ C\\ D\end{array}\begin{pmatrix}0&0&1&1&1\\ 0&0&1&1&1\\ 1&1&0&0&0\\ 1&1&0&0&0\\ 1&1&0&0&0\\ \end{pmatrix}$ (1)
Given the sights of the example above, we define the general GARG-AML method, starting with the adjacency matrix. The first row of the adjacency matrix, which corresponds to entries $(A_{1,j})$, represents the connections of node $v$ under consideration. The next m1m-1 rows, i.e., $(A_{i,j})$ for $2\leq i\leq m$}, represent the connections of nodes that are located at a distance of two from $v$. The final $n$ rows, i.e., $(A_{i,j})$ for $m+1\leq i\leq m+n$}, represent the connections of nodes at a distance of one from $v$.
Enforcing this node ordering—the target node, second-order neighbours, and first-order neighbours—for the adjacency matrix ensures that the diagonal blocks remain sparse (low density), and the off-diagonal blocks have a high density in case we are faced with a pure smurfing pattern. The exact order of nodes within each group is not important. For example, any permutation within the first-order neighbours, such as $C,B,D$, is equally valid.
We consider three primary blocks of the adjacency matrix for score calculation: the upper-left, upper-right, and lower-right blocks. These are defined as follows:
$\text{block}_{1}$ $=$ $(A_{i,j})_{1\leq i,j\leq m}$ (2)
$\text{block}_{2}$ $=$ $(A_{i,j})_{1\leq i\leq m,m+1\leq j\leq m+n}$ (3)
$\text{block}_{3}$ $=$ $(A_{i,j})_{m+1\leq i,j\leq m+n}$ (4)
The density, i.e., the relative number of non-zero elements in each block are essential for our method. GARG-AML measures the similarity between an observed pattern and the pure smurfing template by analysing the expected density and actual density (or sparsity) of each block. This results in a single quantitative score expressing similarity to the pure smurfing pattern.
Several structural properties of the adjacency matrix are taken into account to come to more rigorous density calculations. First, self-loops are excluded; thus, diagonal elements are always zero and are not considered from the sparsity/density calculations. Furthermore, in $\text{block}_{1}$ (upper-left block), the first row and column can never contain ones, as a ‘1’ would indicate a direct connection between a node and its neighbours at distance two—a contradiction. This upper-left block serves as a check that the nodes at distance two do no directly make transactions to each other, in case multiple nodes at a distance of two are present. Similarly, for $\text{block}_{2}$ (upper-right block), all the elements in the first row must be ones since these represent connections between the node under investigation and its first-order neighbours.
We consider only the free entries, i.e., those that are not certain upfront, for density calculations. For readability, we define the block size as the number of free elements. This is reflected in both the numerator (number of non-zero elements) and denominator (block size).
$\text{score}_{1}$ $=$ $\frac{\text{sum}(\text{block}_{1})}{m^{2}-3m+2}$ (5)
$\text{score}_{2}$ $=$ $\frac{\text{sum}(\text{block}_{2})-m}{mn-n}$ (6)
$\text{score}_{3}$ $=$ $\frac{\text{sum}(\text{block}_{3})}{n^{2}-n}$ (7)
Smurfing is likely to be present when the density of $\text{block}_{2}$ (upper-right block) is high, while $\text{block}_{1}$ (upper-left block) and $\text{block}_{3}$ (lower-right) exhibit densities close to zero. Therefore, the final GARG-AML score is calculated as=:
$\text{score}_{\text{total}}=\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}},$ (8)
where $l_{i}$ is a weight given to each term in the score calculation. The inclusion of $l_{i}$ increases the flexibility of the score calculation. If $l_{i}=1$, the score is the average of $\text{score}_{1}$ and $\text{score}_{3}$. In our experiments, however, we take a weighted average using the block sizes:
$l_{1}$ $=$ $m^{2}-3m+2$
$l_{3}$ $=$ $n^{2}-n.$
When there is only one destination account, there is only one node at distance two (as in Figure 2). This results in $l_{1}$ being 0. Then, the calculations simplify to $\text{score}_{\text{total}}=\text{score}_{2}-\text{score}_{3}$, and we only analyse whether all money ends up at the destination node (high $\text{score}_{2}$) and none of the intermediary accounts interact with each other (low $\text{score}_{3}$).
The final score in Equation (8) ranges from 1-1 to $1$. When the second-order neighbourhood resembles a smurfing pattern, the final score is close to 1. To illustrate how these calculations are done, additional examples are provided in Appendix A.1.
3.3 Directed
A defining trait of smurfing is the uni-directional flow of funds. Money moves from one account to others with minimal or no return flow (Starnini et al., 2021; Lee et al., 2020; Li et al., 2020; Shadrooh & Nørvåg, 2024), forming what we call pure smurfing patterns (see Figure 3). To capture this behaviour, we extend GARG-AML to directed networks.
Figure 3: The directed version of the smurfing pattern example, extended with an additional sending node ($A^{\prime}$) and receiving node ($E^{\prime}$). Nodes at level 0 are put in blue, at level 1 in purple and at level 2 in green.
We, again, start with an example. In the directed network of Figure 3, unlike the undirected case (Figure 2), node $A$ only sends funds to the smurfs, while node $E$ only receives from them. This leads to a non-symmetric adjacency matrix, requiring the analysis of nine matrix blocks instead of three (see Equation (9)). To ensure consistency in block-wise density calculations—especially with multiple senders or receivers—nodes at distance two are classified as senders (level 0) and receivers (level 2). When considering the network on the right in Figure 3, we want that node $A^{\prime}$ is put at the same level as node $A$ and node $E^{\prime}$ at the same level as node $E$. The resulting adjacency matrix is then
$\begin{array}[]{l}A\\ A^{\prime}\\ \hline\cr B\\ C\\ D\\ \hline\cr E\\ E^{\prime}\end{array}\left(\begin{array}[]{cc|ccc|cc}0&0&1&1&1&0&0\\ 0&0&1&1&1&0&0\\ \hline\cr 0&0&0&0&0&1&1\\ 0&0&0&0&0&1&1\\ 0&0&0&0&0&1&1\\ \hline\cr 0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0\end{array}\right).$ (9)
Accordingly, we define two types of second-order neighbourhoods in network $G(V,E)$ for the general GARG-AML score calculations. The strong second-order neighbourhood, denoted $\mathcal{N}^{s}_{2}(v)$, considers nodes reachable from $v$ via a directed path of length two. The weak second-order neighbourhood, $\mathcal{N}^{w}_{2}(v)$, is based on the undirected network. In both, only nodes at exact distance two are included.
A node $w$ is assigned to level 0 if
$w\in\mathcal{N}^{w}_{2}(v)\setminus\mathcal{N}^{s}_{2}(v).$ (10)
To account for reverse flow patterns we repeat this selection on the reversed network. This means that a node $w$ a distance two of node $v$ is put at level 0 if there is no directed path of length two from $v$ to $w$ nor from $w$ to $v$.
As in the undirected case, structural constraints of the adjacency matrix are considered when calculating block-wise densities. For the upper-left block, $\text{score}_{00}$, the first row and column are excluded since all other level-0 nodes are nodes at distance two. Additionally, diagonal elements in all on-diagonal blocks ($\text{score}_{00}$, $\text{score}_{11}$ and $\text{score}_{22}$) are excluded from density calculations, as self-loops are not present.
Adjustments are also made to $\text{score}_{20}$ and $\text{score}_{02}$, which represent connections between the node and its second-order neighbours. The first row of $\text{block}_{02}$ (from the node to its second-order neighbours) and first column of $\text{block}_{20}$ (from the second-order neighbours to the node) are always zero and are therefore excluded.
No adjustments are applied to the remaining blocks, as the directionality among level-1 and level-2 nodes does not impose clear structural constraints.
We take the sizes to be $\mid\mathcal{N}_{1}(v)\mid=n$, $\mid\mathcal{N}^{s}_{2}(v)\mid=m$, and $\mid\mathcal{N}^{w}_{2}(v)\setminus\mathcal{N}^{s}_{2}(v)\mid=l-1$, resulting in $l$ nodes at level 0, $n$ nodes at level 1, and $m$ nodes at level 2. The full adjacency matrix is partitioned into nine blocks, as shown in Equation (11), with block labels using zero-based indexing to reflect node levels 0, 1, and 2.
$\left(\begin{array}[]{ccc}\text{b}_{00}&\text{b}_{01}&\text{b}_{02}\\ \text{b}_{10}&\text{b}_{11}&\text{b}_{12}\\ \text{b}_{20}&\text{b}_{21}&\text{b}_{22}\end{array}\right),$} (11)
where
$\text{b}_{00}$ $=$ $(A_{i,j})_{1\leq i,j\leq l}$
$\text{b}_{01}$ $=$ $(A_{i,j})_{1\leq i\leq l,l+1\leq j\leq l+n}$
$\text{b}_{02}$ $=$ $(A_{i,j})_{1\leq i\leq l,l+n+1\leq j\leq l+n+m}$
$\text{b}_{10}$ $=$ $(A_{i,j})_{l+1\leq i\leq l+n,1\leq j\leq l}$
$\text{b}_{11}$ $=$ $(A_{i,j})_{l+1\leq i,j\leq l+n}$
$\text{b}_{12}$ $=$ $(A_{i,j})_{l+1\leq i\leq l+n,l+n+1\leq j\leq l+n+m}$
$\text{b}_{20}$ $=$ $(A_{i,j})_{l+n+1\leq i\leq l+n+m,1\leq j\leq l}$
$\text{b}_{21}$ $=$ $(A_{i,j})_{l+n+1\leq i\leq l+n+m,l+1\leq j\leq l+n}$
$\text{b}_{22}$ $=$ $(A_{i,j})_{l+n+1\leq i,j\leq l+n+m}$
To compute the final score, we take the weighted average of the two blocks expected to have high density — according to the smurfing pattern in Equation (9) — and subtract the weighted average of the remaining blocks, which are expected to be sparse. The use of the average ensures that the final GARG-AML score is between -1 and 1, to have consistent comparison and interpretation across datasets.
score $=$ $\text{mean}(\text{score}_{01},\text{score}_{12})$
$-\text{mean}(\text{score}_{00},\text{score}_{02},\text{score}_{10},\text{score}_{11},\text{score}_{20},\text{score}_{21},\text{score}_{22})$
To illustrate how these calculations are done, a small example is provided in Appendix A.2.
The GARG-AML scores offer three key advantages. First, the scores are close to 1 when a node’s second-order neighbourhood resembles a smurfing pattern, providing a clear and interpretable signal for both automated detection and human investigation. Second, the scores can easy be integrated into existing workflows. The densities and final score can be seen as additional feature values that can be added when training a model or setting up business rules. Third, the method is highly scalable. Each score is computed independently using only a node’s second-order neighbourhood, allowing for efficient parallel computation and low memory usage, even on large transaction networks.
4 Experimental Setup
4.1 Data
We experimentally evaluate the detection power of GARG-AML on multiple datasets. A first set concerns synthetic networks that we generate ourselves using various existing methods, to illustrate how GARG-AML scales with growing transaction networks and to evaluate how well it operates when faced with camouflaging tactics. In addition, the performance of GARG-AML is tested on two realistic open-source datasets. These mimic real-life transaction data; they include millions of observations of transactions, with only very few of these confirmed money laundering cases.
4.1.1 Synthetic Data
We generate synthetic networks of varying sizes using three generation methods, i.e., Barabási-Albert (Barabási & Albert, 1999), Erdős-Rényi (Erdős & Rényi, 1959) and Watts–Strogatz (Watts & Strogatz, 1998). The resulting networks have a range of different properties, e.g., scale-free, small world, etc., that allow the evaluation of GARG-AML across different network topologies.
Each of these methods has specific parameters that we vary to obtain a range of networks to test our method across different conditions. We summarize the ranges of these parameters in Table 3. The first parameter is the size of the network, which is subsequently set to 100, 10,000 and 100,000.
Table 3: Overview of the parameters used to construct the synthetic data.
Parameter Values Barabási-Albert Erdős-Rényi Watts–Strogatz
Number of initial nodes {100; 10,000;100,000}
Number of edges { 1; 2; 5 }
Edge (rewiring) probability { 0.001; 0.01 }
Number of smurfing patterns { 3; 5 }
Total number of combinations 18 12 36
The other parameters have slightly different definitions depending on the method. The number of edges is set at 1, 2 or 5. For Barabási-Albert, this represents the number of nodes to which each new node connects, resulting in a large difference in degree across the nodes. For Watts–Strogatz, on the other hand, this is half the initial degree that all the nodes have before some of the edges are rewired.
The edge (rewiring) probability is set to 0.001 and 0.01. The Erdős-Rényi method uses this parameter to express the probability of having an edge between any two nodes. The Watts–Strogatz method starts with a fixed number of nodes, and uses this parameter to express the probability that an edge changes its end node (rewiring).
Smurfing patterns are injected after network generation. The three types of smurfing patterns that we consider in the experiments, and which are visualized in Figure 4, are the following:
  1. 1.
    Separate: The smurfing patterns are not connected to the rest of the network;
  2. 2.
    External smurfs: Mules are added between two existing entities in the network;
  3. 3.
    Internal smurfs: Existing nodes in the network are assigned the roles of sender, receiver and mule, and corresponding transactions are added.
Figure 4: Illustration of a random network in which three different methods of injecting smurfing patterns are presented. The solid lines indicate nodes and edges of the original network, whereas the dotted lines indicate injected nodes and edges that correspond to smurfing.
In our experiments, each of the smurfing patterns is included either three or five times; see Table 3. The number of smurfs in each pattern is randomly chosen between two and ten. This is done to allow for smaller as well as larger smurfing patterns. A summary of all the synthetic datasets obtained is presented in Table 4.
Table 4: Summary of the numbers of nodes and edges and money laundering percentages for the synthetic datasets according to the considered network generation parameters.
Dataset n_nodes m_edges p_edges n_patterns Nodes Edges Labels
Barabási-Albert 100 1 - 3 145 199 46.9 $\%$
Barabási-Albert 100 2 - 3 155 345 60.0 $\%$
Barabási-Albert 100 5 - 3 141 582 50.35 $\%$
Erdős-Rényi 100 - 0.001 3 151 122 52.32 $\%$
Erdős-Rényi 100 - 0.01 3 153 155 60.78 $\%$
Watts-Strogatz 100 1 0.001 3 144 204 48.61 $\%$
Watts-Strogatz 100 1 0.01 3 146 216 52.05 $\%$
Watts-Strogatz 100 2 0.001 3 131 268 40.46 $\%$
Watts-Strogatz 100 2 0.01 3 146 316 52.74 $\%$
Watts-Strogatz 100 5 0.001 3 151 618 52.32 $\%$
Watts-Strogatz 100 5 0.01 3 145 616 53.1 $\%$
Barabási-Albert 100 1 - 5 180 291 70.0 $\%$
Barabási-Albert 100 2 - 5 179 385 69.83 $\%$
Barabási-Albert 100 5 - 5 175 678 74.29 $\%$
Erdős-Rényi 100 - 0.001 5 182 196 70.33 $\%$
Erdős-Rényi 100 - 0.01 5 163 161 65.03 $\%$
Watts-Strogatz 100 1 0.001 5 176 271 65.91 $\%$
Watts-Strogatz 100 1 0.01 5 152 248 69.74 $\%$
Watts-Strogatz 100 2 0.001 5 171 376 70.18 $\%$
Watts-Strogatz 100 2 0.01 5 180 380 67.22 $\%$
Watts-Strogatz 100 5 0.001 5 171 669 69.01 $\%$
Watts-Strogatz 100 5 0.01 5 165 655 66.67 $\%$
Barabási-Albert 10,000 1 - 3 10,037 10,101 0.69 $\%$
Barabási-Albert 10,000 2 - 3 10,042 20,105 0.72 $\%$
Barabási-Albert 10,000 5 - 3 10,043 50,103 0.77 $\%$
Erdős-Rényi 10,000 - 0.001 3 10,044 5101 0.71 $\%$
Erdős-Rényi 10,000 - 0.01 3 10,050 50,388 0.79 $\%$
Watts-Strogatz 10,000 1 0.001 3 10,054 10,132 0.84 $\%$
Watts-Strogatz 10,000 1 0.01 3 10,052 10,134 0.85 $\%$
Watts-Strogatz 10,000 2 0.001 3 10,038 20,104 0.7 $\%$
Watts-Strogatz 10,000 2 0.01 3 10,046 20,134 0.85 $\%$
Watts-Strogatz 10,000 5 0.001 3 10,039 50,090 0.63 $\%$
Watts-Strogatz 10,000 5 0.01 3 10,040 50,108 0.72 $\%$
Barabási-Albert 10,000 1 - 5 10,059 10,171 1.15 $\%$
Barabási-Albert 10,000 2 - 5 10,081 20,215 1.38 $\%$
Barabási-Albert 10,000 5 - 5 10,072 50,167 1.2 $\%$
Erdős-Rényi 10,000 - 0.001 5 10,079 5263 1.4 $\%$
Erdős-Rényi 10,000 - 0.01 5 10,067 49,998 1.08 $\%$
Watts-Strogatz 10,000 1 0.001 5 10,081 10,168 1.13 $\%$
Watts-Strogatz 10,000 1 0.01 5 10,070 10,182 1.2 $\%$
Watts-Strogatz 10,000 2 0.001 5 10,075 20,180 1.19 $\%$
Watts-Strogatz 10,000 2 0.01 5 10,071 20,190 1.24 $\%$
Watts-Strogatz 10,000 5 0.001 5 10,081 50,196 1.27 $\%$
Watts-Strogatz 10,000 5 0.01 5 10,060 50,148 1.03 $\%$
Barabási-Albert 100,000 1 - 3 100,044 100,099 0.07 $\%$
Barabási-Albert 100,000 2 - 3 100,038 200,095 0.07 $\%$
Barabási-Albert 100,000 5 - 3 100,043 500,093 0.07 $\%$
Erdős-Rényi 100,000 - 0.001 3 100,033 498,969 0.06 $\%$
Erdős-Rényi 100,000 - 0.01 3 100,044 4,998,019 0.07 $\%$
Watts-Strogatz 100,000 1 0.001 3 100,050 100,124 0.08 $\%$
Watts-Strogatz 100,000 1 0.01 3 100,045 100,096 0.07 $\%$
Watts-Strogatz 100,000 2 0.001 3 100,044 200,114 0.07 $\%$
Watts-Strogatz 100,000 2 0.01 3 100,041 200,116 0.08 $\%$
Watts-Strogatz 100,000 5 0.001 3 100,042 500,114 0.07 $\%$
Watts-Strogatz 100,000 5 0.01 3 100,045 500,116 0.08 $\%$
Barabási-Albert 100,000 1 - 5 100,063 100,167 0.11 $\%$
Barabási-Albert 100,000 2 - 5 100,069 200,175 0.12 $\%$
Barabási-Albert 100,000 5 - 5 100,065 500,159 0.12 $\%$
Erdős-Rényi 100,000 - 0.001 5 100,087 501,039 0.14 $\%$
Erdős-Rényi 100,000 - 0.01 5 100,076 4,997,365 0.13 $\%$
Watts-Strogatz 100,000 1 0.001 5 100,061 100,166 0.11 $\%$
Watts-Strogatz 100,000 1 0.01 5 100,058 100,134 0.1 $\%$
Watts-Strogatz 100,000 2 0.001 5 100,065 200,154 0.11 $\%$
Watts-Strogatz 100,000 2 0.01 5 100,059 200,154 0.11 $\%$
Watts-Strogatz 100,000 5 0.001 5 100,074 500,176 0.12 $\%$
Watts-Strogatz 100,000 5 0.01 5 100,068 500,166 0.11 $\%$
Since we have information on which nodes are involved in which pattern for each dataset, we not only test the methods for money laundering detection in general; we also analyse how the methods perform on separate patterns. Similarly, we investigate how performance changes when money launderers apply higher levels of obfuscation methods.
4.1.2 Open-Source Data
As post-hoc injection of specific patterns may lead to detection bias (Deprez et al., 2024) and altered topology (Egressy et al., 2024), we also include data without this injection in the experiments to evaluate GARG-AML.
We use the HI-Small and LI-Large datasets, made publicly available by Altman et al. (2023) and Egressy et al. (2024), representing a relatively small dataset with high money laundering activity and a large dataset with low money laundering activity, respectively. Table 5 summarises the number of nodes, edges and the percentage of cases labelled as money laundering. Especially LI-Large closely resembles real-world transaction networks.
Table 5: The summary of the two IBM datasets, representing the number of nodes, edges and the percentage of instances labelled as money laundering (ML).
Dataset # nodes # edges percentage ML
HI-Small 515 080 5 078 345 $0.1\%$
LI-Large 2 054 390 176 066 557 $0.05\%$
The two datasets are generated using the same simulator (Altman et al., 2023; Egressy et al., 2024), which models financial transactions in a virtual world involving individuals, companies, and banks. A subset of these transactions is labelled as money laundering. Since the full transaction graph is available, the authors were able to assign detailed labels, categorizing transaction streams into eight distinct types: fan-out, fan-in, gather-scatter, scatter-gather, simple cycle, random, bipartite, and stack.
Among these, the scatter-gather and gather-scatter patterns—illustrated in Figure 1—represent the primary smurfing behaviour that GARG-AML aims to detect. The distribution of labelled patterns is shown in Table 6. It is worth noting that some transactions are labelled as money laundering but do not match any of the eight defined structural patterns. These are grouped under the category “not classified”, indicating that they are suspicious but not attributable to a specific laundering archetype.
Table 6: The distribution of the different types of money laundering transaction patterns for the two synthetic datasets.
Type of pattern HI-Small LI-Large
Fan-out 342 1 985
Fan-in 318 1 935
Gather-scatter 716 3 857
Scatter-Gather 626 3 694
Cycle 287 1 871
Random 191 1 458
Bipartite 263 1 807
Stack 466 2 854
Not classified 1 968 81 143
Total number of transactions 5 078 345 176 066 557
The original labels in these datasets are provided at the edge level, while GARG-AML produces features at the node level. To bridge this gap, we aggregate labels per client by computing the ratio of labelled laundering transactions to the total number of transactions associated with each node. This ratio is interpreted as a propensity score for money laundering.
To generate binary labels, we compare each node’s propensity score against a set of predefined thresholds: $10\%,20\%,30\%,50\%$ and $90\%$. A node is labelled as involved in money laundering if the proportion of its laundering-labelled transactions exceeds the selected threshold. The resulting label distributions are presented in Table 7 and Table 8, corresponding to the HI-Small and LI-Large datasets, respectively.
In many cases, fewer than $0.1\%$ of accounts are assigned a positive label, highlighting the extreme class imbalance and difficulty of the classification task. Additionally, accounts with more than $90\%$ of laundering-labelled transactions are extremely rare, which reflects the realistic behaviour of laundering entities mixing illicit flows with legitimate transactions to obfuscate detection.
Table 7: Percentage of instances labelled as 1 in the HI-Small dataset, for the different patterns at different thresholds.
In percentage $(\%)$ 10 20 30 50 90
Is Laundering 0.445 0.175 0.103 0.020 0.013
Fan-out 0.029 0.012 0.009 0.002 0.002
Fan-in 0.022 0.006 0.005 0.002 0.001
Gather-scatter 0.055 0.018 0.014 0.002 0.001
Scatter-gather 0.036 0.022 0.010 0.003 0.000
Cycle 0.019 0.013 0.005 0.000 0.000
Random 0.014 0.007 0.002 0.000 0.000
Bipartite 0.030 0.007 0.004 0.001 0.001
Stack 0.044 0.021 0.011 0.001 0.001
Table 8: Percentage ($\%$) of instances labelled as 1 in the LI-Large dataset, for the different patterns at different thresholds.
In percentage $(\%)$ 10 20 30 50 90
Is Laundering 0.2162 0.0414 0.0192 0.0017 0.0004
Fan-out 0.0051 0.0020 0.0007 0.0000 0.0000
Fan-in 0.0049 0.0025 0.0015 0.0003 0.0000
Gather-scatter 0.0081 0.0040 0.0029 0.0003 0.0000
Scatter-gather 0.0147 0.0047 0.0022 0.0001 0.0000
Cycle 0.0087 0.0016 0.0006 0.0000 0.0000
Random 0.0056 0.0007 0.0001 0.0000 0.0000
Bipartite 0.0042 0.0010 0.0004 0.0000 0.0000
Stack 0.0095 0.0012 0.0003 0.0000 0.0000
4.2 Model Specifications
The full GARG-AML calculation pipeline consists of multiple stages, beginning with pre-processing steps that prepare the network data for analysis, followed by the computation of the GARG-AML scores. We also introduce a possible extension of the GARG-AML method that applies tree-based methods. We describe this use cases in detail below to illustrate the versatility of GARG-AML. An overview of the complete pipeline is shown in Figure 5 and summarised in Algorithm 1.
Figure 5: The pipeline of the GARG-AML method.
Algorithm 1 GARG-AML Score Calculation
Input: Network $G(V,E)$
Output: GARG-AML propensity score(s)
1:Calculate communities using Louvain method (Blondel et al., 2008)
2:Remove inter-community edges
3:if Directed score then
4:  for node $v\in V$ do
5:   Extract second order neighbourhood of $v$
6:   Distribute neighbours over level 0, 1 and 2
7:   Define the adjacency matrix with level 0, 1 2.
8:   Define the nine blocks using Equation (11)
9:   Calculate score via Equation (3.3)
10:  end for
11:else
12:  for node $v\in V$ do
13:   Extract second order neighbourhood of $v$
14:   Define the adjacency matrix using $v$, $\mathcal{N}_{2}(v)$ and $\mathcal{N}_{1}(v)$
15:   Define the three blocks using Equations (2)-(4)
16:   Calculate score via Equation (8)
17:  end for
18:end if
19:return: scores
Transaction networks often exhibit a scale-free structure (Starnini et al., 2021), which are characterized by the presence of a few nodes (accounts) with disproportionately high numbers of incoming and outgoing transactions. As a result, second-order neighbourhoods can become very large and can encompass an unmanageable number of nodes for score computation.
Therefore, we first partition the network into communities using the Louvain method (Blondel et al., 2008). This method consistently shows state-of-the-art performance for community detection, and scales well to large graphs (Lancichinetti & Fortunato, 2009; Yang et al., 2016). Communities are identified on the basis of the undirected version of the transaction network, and only intra-community edges are retained—edges between communities are discarded.
However, the Louvain method can still produce overly large communities (Fortunato & Barthélemy, 2007; Newman & Girvan, 2004). To mitigate this, we use the implementation from NetworkX (Hagberg et al., 2008), which includes a resolution parameter inspired by the improvements proposed by Traag et al. (2019). In our experiments, we set this parameter to 10, which effectively encourages the algorithm to form smaller, more manageable communities, thus maintaining tractability in the GARG-AML score computation.
After pre-processing the network using the Louvain method, we calculate the block densities as described above via Equation (8) and Equation (3.3).
As a possible extension, we apply a decision tree and gradient boosting classifier with features on the basis of the GARG-AML scores. These include the original score as before, which is enhanced with summary statistics of each node’s first-order neighbours, i.e., the minimal, mean and maximal values of the GARG-AML scores as well as the standard deviation of the neighbours’ scores. Since successful smurfing involves many smurfs, we also include each node’s degree and the summary statistics—minimum, mean, maximum and standard deviation—of the degrees of the node’s neighbours.
The goal of including these extra features is to capture more complex neighbourhood structures that may not be evident from the scores alone. In particular, this helps reduce false positives—for example, a linear chain of three nodes may yield a perfect score, even though a single “smurf” is insufficient evidence of an actual laundering operation.
Applying this on the undirected and directed network results in four additional model. We use a 70-30 stratified train-test split for evaluation. This split is performed on the resulting feature table, not the network itself. For the decision tree, we use the Gini criterion and require that each leaf contain at least ten samples. The gradient boosting classifier consists of 100 trees with a maximal depth of three and a learning rate of 0.1. We require that each leaf node should contain at least 10 observations.
We benchmark GARG-AML and its extension against established state-of-the-art methods, namely, AutoAudit (Lee et al., 2020) and FlowScope (Li et al., 2020). For both methods, we use the implementations provided by the authors and retain their default settings.
Various small adaptations are necessary. AutoAudit returns an optimal reordering of indices in the adjacency matrix. However, this reordering affects only part of the matrix and may depend on the arbitrary labelling of nodes. To ensure robustness, we repeat the experiments three times with random initialization of the node indices and report the average results. Both methods also return an ordering of transactions resulting from the $n$ different iterations. To reflect the decreasing level of certainty, the first group gets prediction $1$, the second $\frac{n-1}{n}$ and so on until the final, $n$th, group gets score $\frac{1}{n}$. All other nodes that were not returned by the methods get score 0.
4.3 Evaluation
Two key aspects are critical for business adoption. The first is model performance—a robust model must accurately detect money laundering while keeping false positives to a minimum. The second is scalability. Real-world transaction datasets are massive, with banks processing thousands to millions of transactions daily. Therefore, any proposed method must scale efficiently to such volumes. Accordingly, the evaluation of the methods focuses on these two dimensions.
For performance evaluation, the GARG-AML scores are assessed against ground-truth labels. The evaluation is conducted using the area under the receiver operating characteristic curve (AUC-ROC) and the area under the precision-recall curve (AUC-PR). These scores are also used to rank the methods, where the best-performing method receives rank 1, the second-best method receives rank 2, and so on.
While AUC-ROC is a widely used metric for binary classification, it may not be sufficient for anti-money laundering (AML) datasets, which are often highly imbalanced—typically with fewer than $0.1\%$ observations flagged as laundering money. To address this imbalance, we include AUC-PR, which has been shown to be a more reliable performance metric under situations of extreme label imbalance (Davis & Goadrich, 2006; Ozenne et al., 2015). In general terms, the AUC-PR is the expected average precision when a random recall threshold is selected. In other words, the AUC-PR measures the model’s ability to identify positive examples correctly while minimising false positives across all possible threshold values. The AUC-PR specifically focuses on the model’s precision in identifying the positive class at various recall thresholds.
Scalability is assessed in terms of calculation time. In addition to runtime, memory usage is also a key factor. We also report any instances where methods fail due to memory limitations. The theoretical time complexities are $O\left(k\mid E\mid\log\left(\mid V\mid\right)\right)$ for Flowscope (Li et al., 2020) and $O\left(\mid V\mid\cdot\mid E\mid+\mid V\mid^{2}\right)$ for AutoAudit (Lee et al., 2020).
For GARG-AML, the main calculations for each individual node are the selection of the second-order neighbourhood and the assignment of the neighbours to the levels (in the directed case). Assuming that we are working in sparse graphs, both steps have complexity $O\left(\mid V\mid\right)$. This brings us to a total worst case time complexity of $O\left(n^{2}\right)$. Hence, theoretically, GARG-AML scales as AutoAudit, while Flowscope is much faster. However, we can speed up the calculations by doing running them in parallel. In addition, second-order neighbours are generally much smaller than $n$, meaning that also those steps can run very quickly.
All the experiments were conducted on the Flemish Supercomputer (VSC) with an Intel Xeon Platinum 8360Y CPU and 200 GB of memory. Each experiment was constrained to a time budget of 16 hours, the maximum on the VSC.
4.4 Statistical Test
The statistical significance of the difference in ranks is evaluated using the Friedman test (Friedman, 1937, 1940) and the post hoc Nemenyi test (Nemenyi, 1963), as described by Demšar (2006). It compares the performance of the $k$ different methods across the $N$ datasets. If we take $r_{i}^{j}$ to be the rank of method $j$ on dataset $i$ for some performance metric, then the average performance across all datasets is equal to $R_{j}=\frac{1}{N}\sum_{i}r_{i}^{j}$. In our experiments, the Friedman and Nemenyi statistical tests compare the ranks of the 8 $(k)$ methods across the 66 $(N)$ synthetic datasets. We apply a significance level of $0.05$ for these tests.
The null hypothesis of the Friedman test is that the average ranks $R_{j}$ are equal, and the Friedman statistic is equal to
$Q=\frac{12N}{k(k+1)}\left[\sum_{j}R_{j}^{2}-\frac{k(k+1)^{2}}{4}\right].$
The Friedman statistic follows a $\chi^{2}_{k-1}$ distribution under this null hypothesis.
Rejecting the null hypothesis indicates that significant differences are present among the methods. We then proceed with the post hoc Nemenyi test, which finds that the performances of two methods are significantly different at level $\alpha$ if their respective average ranks differ by at least the critical difference (CD):
$\text{CD}=q_{\alpha}\sqrt{\frac{k(k+1)}{6N}},$
where critical levels $q_{\alpha}$ are provided by Demšar (2006). We visualize the results using CD diagrams (Demšar, 2006) for the AUC-ROC and AUC-PR.
5 Results and Discussion
The results are presented in two parts. First, we present and discuss the scalability of the methods. Afterward, we show the results for the performance and the CD diagrams.
5.1 Scalability
Figure 7 and Figure 7 hows the calculation times across the various network sizes of the synthetic data and IBM datasets, respectively. The time in Figure 7 is shown on a logarithmic scale since this provides a better distinction among the methods across dataset sizes. These results indicate that FlowScope is the fastest of all methods, although GARG-AML seems to also scale well to larger datasets, contrary to what we expected from the theoretical worst-case time complexity. In contrast, AutoAudit—although quite fast in the beginning—has difficulties with networks of the largest size. In addition, the distribution of the calculation time of AutoAudit is much wider and has more outliers, which indicates that it is highly sensitive to the network topology.
Figure 6: Box plot of the run time for each method on a logarithmic scale over the three different sizes of synthetic datasets. From left to right are FlowScope, AutoAudit, undirected GARG-AML and directed GARG-AML scores.
Figure 7: Bar plot of the run time for each method on a logarithmic scale over the two IBM datasets. From left to right are FlowScope, undirected GARG-AML and directed GARG-AML scores.
Another key observation is that AutoAudit consistently runs out of time—exceeding 16 hours—on Watts-Strogatz networks of size 10,000. This is reflected by setting the out-of-time results equal to 16 hours (57,600 seconds) in Figure 7, which results in outliers for AutoAudit. For the larger networks, i.e. the synthetic network of size 100,000 and the IBM datasets, AutoAudit exceeds the available memory. Hence, the results for AutoAudit are omitted entirely from Figure 7.
5.2 Performance
5.2.1 Synthetic Data
The results are presented according to the different injected patterns. These are provided in Figure 8, with their corresponding ranks presented in Figure 9.
Figure 8: AUC-ROC and AUC-PR values of the compared methods across all synthetic datasets, with the split-on money laundering pattern.
Figure 9: Ranks of the AUC-ROC and AUC-PR values of the compared methods across all synthetic datasets, with the split-on money laundering pattern, where rank 1 indicates the best performance and rank 8 indicates the worst performance.
The AUC-PR values are more spread out than the AUC-ROC values. This is caused by the varying level of class imbalance, as shown in Table 4. For a model with random label predictions, the AUC-ROC is always 0.5, regardless of the label distribution, whereas the AUC-PR equals the proportion of positive labels, which leads to the high variance shown in Figure 8.
Figure 8 indicates that, on average, the undirected GARG-AML scores consistently showing great performance. This performance is boosted when combining the GARG-AML scores with the tree-based learners. In comparison, the performance of AutoAudit is less stable. Here, we set all the AUC-ROC and AUC-PR values to 0 when the method runs out of time or out of memory to allow for a full comparison of the ranks. For these cases, AutoAudit ranks last compared with the other methods.
The ranks in Figure 9 corroborate the results from before. We can see more clearly that combining the tree-based learners with the scores and summary statistics of the node’s neighbours score better than the other methods, even when just applying a simple decision tree. We test the statistical significance of the difference in average rank using the Friedman test. For all the cases, the Friedman test rejects the null hypothesis, which indicates that significant differences between the ranks are present.
The post hoc Nemenyi test is used to discern which methods have significantly different average ranks. The results are shown in Figure 11 and Figure 11 for the AUC-ROC and AUC-PR, respectively. Here, again the tree-based methods differ significantly from most other methods, and the undirected scores on their own also obtain strong performance. For the existing mules patterns, however, it seems that the best performing models do not differ significantly from FlowScope.
Figure 10: Critical distance plots of the average ranks according to the AUC-ROC for the considered patterns.
Figure 11: Critical distance plots of the average ranks according to the AUC-PR for the considered patterns.
5.2.2 IBM Data
This section presents the results on the IBM datasets for the smurfing patterns, i.e., gather-scatter and scatter-gather, and the full ‘Is Laundering’ label. The thresholds applied for binary label definition are 0.1, 0.2 and 0.9. The results for the HI-Small and LI-Large datasets are presented in Figure 12 and Figure 13, respectively. The results for all patterns and thresholds are provided in Appendix B.
Figure 12: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset on the money laundering and smurfing labels across different thresholds.
Figure 13: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset on the money laundering and smurfing labels across different thresholds. Some panels are empty, due to the absence of positive labels for these cases.
Similar to the results on the synthetic data, the undirected version of the GARG-AML scores obtain best performance for almost all cases for the AUC-ROC values, followed by the tree-based models. For both datasets, the undirected GARG-AML scores give great results for the general laundering cases as well, which include more patterns than just smurfing.
For the lower threshold values, we see a significant improvement of AUC-PR across the tree-based methods. This results from these methods being able to better capture the interaction among a node’s GARG-AML score and those of its neighbours. This reduces the false positives compared to the basic score, resulting in a much higher AUC-PR. Therefore, on lower thresholds, we conclude that the tree-based methods are better suited for smurfing detection.
For the highest thresholds the basic GARG-AML scores provide great performance, with AUC-ROC above $80\%$. In these cases, only clients are labelled that have $90\%$ or more of their transactions involved in money laundering. Therefore, low levels of obfuscation are present and the transaction patterns are very close to pure smurfing, making them easily recognisable by GARG-AML.
The tree-based models, on the other hand, seem to perform poorly, in line with FlowScope. It seems that the class imbalance (see also Table 7 and Table 8) is too extreme in these cases, resulting in the model’s inability to train properly.
6 Limitations
This work has limitations related to both data availability and methodology.
6.1 Limitations in the Data
GARG-AML assumes full visibility into all transactions, which is rarely the case in practice (Altman et al., 2023; Tariq & Hassani, 2025; Dumitrescu et al., 2022). Banks typically have access only to transactions that involve their own clients, including sender and receiver account numbers and client attributes. Transactions between two external accounts remain hidden, which makes second-order neighbourhoods incomplete. This can obscure smurfing structures or lead to incorrect classification (see Figure 14). Moreover, banks cannot link multiple third-party accounts that are owned by the same individual, thus further weakening pattern recognition.
Figure 14: Illustration of smurfing over different banks, which are represented by different colours. The (i)-icon indicates that the main (blue) bank has all client information about the corresponding account. The dotted lines indicate that the bank is unaware of these transactions/accounts. In most cases, the blue bank is aware of a big part of the transaction network.
This is mainly a limitation when only one bank would be guarding against money laundering. When two banks are involved, one of these two will have a clear picture of the transaction scheme. Although the transactions go through different banks in Figures 14(a) and (b), the blue bank still has the full transaction pattern. It becomes much harder to have a full picture when more than two banks are involved.
Given these limitations, we advocate for stronger collaboration both among banks and across countries. By enabling the secure sharing of transaction information, many of the current blind spots could be substantially reduced. In such an environment, GARG-AML would be positioned to deliver significantly greater value.
6.2 Limitations in the Methodology
GARG-AML uses unweighted graph structures, ignoring transaction volume and timing. While these are simplifications, the resulting smurfing score remains interpretable and useful in downstream AML systems. Timing issues can be addressed by restricting analysis to shorter time windows, applying sliding windows, or comparing scores across temporal blocks.
7 Conclusion
This paper introduces Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), which is a novel approach for detecting smurfing by quantifying the structural similarity of nodes’ neighbourhoods to known money laundering patterns via density-based features in the adjacency matrix. We validated GARG-AML on several synthetic datasets and compared it to state-of-the-art smurfing detection methods. The results of the experiments show that GARG-AML achieves strong detection power, with state-of-the-art performance of GARG-AML scores, and excellent scalability for large networks. Unlike approaches that are based on complex optimization techniques or machine learning models, GARG-AML relies on domain expertise and straightforward density-based calculations.
While GARG-AML is not a complete solution for smurfing detection, it offers two key advantages over existing methods: (1) it produces interpretable features—either as separate densities or a single composite score—that can be seamlessly integrated into AML systems, and (2) it is computationally efficient, with parallelisable and memory-light calculations, thus enabling practical deployment in real-world settings.
Future work will explore (i) treating individual densities as features in classification models and (ii) extending the methodology to incorporate transaction timing and volume, as well as generalising it to other anomalous transaction patterns.
Additionally, money laundering schemes typically exploit cross-bank and cross-border transaction chains, which fragment the available information and obscure detection efforts. This work advocates that enhanced collaboration, both at the national level among domestic banks and at the international level across jurisdictions, is essential. Establishing secure data-sharing frameworks and cooperative monitoring mechanisms would significantly strengthen the ability to trace complex laundering networks and improve the overall effectiveness of anti-money laundering measures.
Finally, detecting smurfs is an important research area for network security and intrusion detection (Yang et al., 2009; Shadrooh & Nørvåg, 2024). Further research could analyse whether similar featurisation of the patterns is possible for denial of service (DoS) or similar attacks.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Toy Example GARG-AML Scores
We demonstrate the calculations and results as described in Section 3, using a toy example. We start with a Watts-Strogatz network containing 20 nodes, each connecting to the next two nodes (initial degree equal to four), with a rewiring probability of $20\%$. We inject a smurfing pattern, where money originates at a new node and gets transferred to an existing node in the network using three mule accounts. The full network is visualised in Figure A.1.
We select two nodes to illustrate how the GARG-AML scores are calculated. We select node 23 because its second-order neighbourhood consitutes a pure smurfing pattern. We also select node 8, since its neighbours interact a lot with each other. So, the calculations for node 8 illustrate how the scores are impacted by these interactions.
Figure A.1: The undirected network with toy example. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
A.1 Undirected Calculations
We start with the pure smurfing pattern using node 23. Its second-order neighbourhood is visualised in Figure A.2. Based on this neighbourhood, we construct the adjacency matrix as follows:
$\begin{array}[]{r}23\\ 19\\ 20\\ 21\\ 22\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&1&1&1\\ {\color[rgb]{0,1,0}1}&1&{\color[rgb]{1,0,0}0}&0&0\\ {\color[rgb]{0,1,0}1}&1&0&{\color[rgb]{1,0,0}0}&0\\ {\color[rgb]{0,1,0}1}&1&0&0&{\color[rgb]{1,0,0}0}\\ \end{pmatrix}$ (13)
The elements that should be zero, and hence are not taken into account, are indicated in red. Those elements that are certainly equal to one are indicated in green.
Figure A.2: Visualisation of the undirected second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
We calculate the scores by taking the upper-left block as block 1, the upper-right as block 2, and the lower-right as block 3. Using the notation as in Equations (5)-(8) , this results in the density calculations and sizes, where the size is the number of free elements:
$\text{score}_{1}$ $=$ $0$
$\text{score}_{2}$ $=$ $1$
$\text{score}_{3}$ $=$ $0$
$l_{1}$ $=$ $0$
$l_{2}$ $=$ $3$
$l_{3}$ $=$ $6$
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}}$ (14)
$=$ $1$
As expected, the score that we assign to node 23 is 1. This indicates that its second-order neighbourhood forms a pure smurfing pattern.
The second-order neighbourhood of node 8 is shown in Figure A.3. We observe an interaction between node 7 and 9, which are part of the neighbours at distance one. Additionally, multiple interactions between nodes at distance two can be observed. This already indicates that the neighbourhood does not represent a pure smurfing pattern. The adjacency matrix is given in Equation (15), with the same colour scheme as before.
Figure A.3: Visualisation of the undirected second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
$\begin{array}[]{r}8\\ 3\\ 4\\ 5\\ 10\\ 11\\ 15\\ 6\\ 7\\ 9\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}&{\color[rgb]{0,1,0}1}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&1&1&0&0&1&0&0&1\\ {\color[rgb]{1,0,0}0}&1&{\color[rgb]{1,0,0}0}&1&1&0&0&0&1&0\\ {\color[rgb]{1,0,0}0}&1&1&{\color[rgb]{1,0,0}0}&0&0&0&0&1&0\\ {\color[rgb]{1,0,0}0}&0&1&0&{\color[rgb]{1,0,0}0}&1&0&1&0&0\\ {\color[rgb]{1,0,0}0}&0&0&0&1&{\color[rgb]{1,0,0}0}&0&1&0&0\\ {\color[rgb]{1,0,0}0}&1&0&0&0&0&{\color[rgb]{1,0,0}0}&0&0&1\\ {\color[rgb]{0,1,0}1}&0&0&0&1&1&0&{\color[rgb]{1,0,0}0}&0&1\\ {\color[rgb]{0,1,0}1}&0&1&1&0&0&0&0&{\color[rgb]{1,0,0}0}&0\\ {\color[rgb]{0,1,0}1}&1&0&0&0&0&1&1&0&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (15)
The first row contains the node we will give a score, here node 8. The following rows represent the nodes at distance two, i.e., 3, 4, 5, 10, 11 and 15. The final rows represent the nodes at distance one, i.e., 6, 7 and 9.
This results in the following scores:
$\text{score}_{1}$ $=$ $0.4$
$\text{score}_{2}$ $=$ $0.33$
$\text{score}_{3}$ $=$ $0.33$
$l_{1}$ $=$ $30$
$l_{2}$ $=$ $18$
$l_{3}$ $=$ $6$
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}}$ (16)
$=$ $-0.0556$
As indicated above, node 8 is not involved in a smurfing scheme. This is corroborated by the fact that the final score is negative.
A.2 Directed Calculations
We demonstrate the calculations of the directed GARG-AML scores on the same network as before, but with directed edges. This network is illustrated in Figure A.4. We again calculate the scores for node 23 and 8.
Figure A.4: The directed network with toy example. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
The directed second-order neighbourhood for node 8 is shown in Figure A.5, and the corresponding adjacency matrix is provided in Equation (17). Note that we now have a different ordering of the nodes, using level 0 (node 23), level 1 (nodes 20, 21, 22) and level 2 (node 19), resulting in nine blocks, as defined in Section 3.3.
$\begin{array}[]{r}23\\ 20\\ 21\\ 22\\ 19\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&1&1&1&{\color[rgb]{1,0,0}0}\\ 0&{\color[rgb]{1,0,0}0}&0&0&1\\ 0&0&{\color[rgb]{1,0,0}0}&0&1\\ 0&0&0&{\color[rgb]{1,0,0}0}&1\\ {\color[rgb]{1,0,0}0}&0&0&0&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (17)
Figure A.5: Visualisation of the directed second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
The density of the nine blocks is calculated. We will use the matrix representation as in Equation (11) to represent the densities of the blocks. For the adjacency matrix in Equation (17), this results in the following nine densities:
$\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 0&0&0\end{array}\right)$} (18)
The score for node 23 is then calculated as
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{mean}(\text{score}_{01},\text{score}_{12})$
$-\text{mean}(\text{score}_{00},\text{score}_{02},\text{score}_{10},\text{score}_{11},\text{score}_{20},\text{score}_{21},\text{score}_{22})$
$=$ $\text{mean}(1,1)-\text{mean}(0,0,0,0,0,0,0)$
$=$ $1.$ (20)
This again shows that the pure smurfing pattern gets a perfect score.
We do the same exercise for node 8. Here, level 0 consists of nodes 8 and 15, level 1 consists of nodes 6, 7 and 9, and level 2 consists of nodes 3, 4, 5, 10 and 11. Contrary to the second-order neighbourhood of node 23, we see on Figure A.6 that money flows in multiple directions, with nodes sending and receiving money to and from nodes across levels. This is already an indication that node 8 is not part of a smurfing scheme. The corresponding adjacency matrix is given in Equation (21).
Figure A.6: Visualisation of the directed second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
$\begin{array}[]{r}8\\ 15\\ 9\\ 6\\ 7\\ 3\\ 4\\ 5\\ 10\\ 11\end{array}\begin{pmatrix}{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&1&0&0&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&0&0&0&0&0&0&0&0\\ 0&0&{\color[rgb]{1,0,0}0}&0&0&0&0&0&1&1\\ 1&0&0&{\color[rgb]{1,0,0}0}&0&0&0&0&0&0\\ 1&1&1&0&{\color[rgb]{1,0,0}0}&0&0&0&0&0\\ {\color[rgb]{1,0,0}0}&1&0&0&1&{\color[rgb]{1,0,0}0}&1&1&0&0\\ {\color[rgb]{1,0,0}0}&0&0&1&0&0&{\color[rgb]{1,0,0}0}&1&1&0\\ {\color[rgb]{1,0,0}0}&0&0&1&0&0&0&{\color[rgb]{1,0,0}0}&0&0\\ {\color[rgb]{1,0,0}0}&0&0&0&0&0&0&0&{\color[rgb]{1,0,0}0}&1\\ {\color[rgb]{1,0,0}0}&0&0&0&0&0&0&0&0&{\color[rgb]{1,0,0}0}\end{pmatrix}$ (21)
For this network, we get the following nine densities:
$\left(\begin{array}[]{ccc}0&0.1667&0\\ 0.5&0.1667&0.133\\ 0.2&0.2&0.25\end{array}\right)$} (22)
We observe that the densities are much different than for the pure smurfing pattern. This translates to the following score calculation:
$\text{score}_{\text{GARG-AML}}$ $=$ $\text{mean}(\text{score}_{01},\text{score}_{12})$
$-\text{mean}(\text{score}_{00},\text{score}_{02},\text{score}_{10},\text{score}_{11},\text{score}_{20},\text{score}_{21},\text{score}_{22})$
$=$ $\text{mean}(0.1667,0.133)-\text{mean}(0,0,0.5,0.1667,0.2,0.2,0.25)$
$=$ $0.15-0.188$ (24)
$=$ $-0.038.$ (25)
The score for node 8 is below 0. Hence, we have no suspicion that this node is part of a smurfing scheme.
Appendix B Full Results IBM Data
In this part of the appendix, we provide the full results on the IBM datasets. These include all different patterns defined by Altman et al. (2023). The binary labels are defined using the threshold values 0.1, 0.2, 0.3, 0.5 and 0.9. The results on the HI-Small and LI-Large datasets are provided in Figure B.1 and Figure B.2, respectively.
Figure B.1: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset across all thresholds and patterns.
Figure B.2: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset across all thresholds and patterns.